find the inverse function. a.) f(x)=6-3x, b.)f(x)= 3 square root of x c.)f(x)= x^2+2
Answer by Cromlix(4381)

Hi there,
Inverse function.
This is how I do them:-
a) f(x) = 6 - 3x
y = 6 - 3x
Change the x for the y and the y for the x:-
x = 6 - 3y
Subtract 6 from both sides
x - 6 = 6 - 6 - 3y
x - 6 = -3y
Multiply both sides by -1
6 - x = 3y
Divide both sides by 3
(6 - x)/3 = 3y/3
y = (6 - x)/3
...................
b) f(x) = 3 square root x
y = 3 square root x
Change x for y and y for x
x = 3 square root y
Divide both sides by 3
x/3 = 3 square root y/3
x/3 = square root y
Square both sides
(x/3)^2 = (square root y)^2
(x/3)^2 = y
y = (x/3)^2
....................
c)f(x) = x^2 + 2
y = x^2 + 2
Change x for y and y for x
x = y^2 + 2
Take 2 off both sides
x - 2 = y^2 + 2 - 2
x - 2 = y^2
Square root both sides
Square root(x - 2) = Square root (y^2)
Square root (x - 2) = y
y = Square root (x - 2)
...........................
Hope this helps:-)